### v1

# Compact superconducting vacuum-gap capacitors with low microwave loss and high mechanical coherence for scalable quantum circuits

Abstract
Vacuum gap capacitors have recently gained considerable attention in superconducting circuit platforms due to their compact design and low dielectric losses in the microwave regime. Their ability to support mechanical vibrational modes makes them ideal candidates for circuit optomechanics. However, precise control of gap size and achieving high coherence in mechanical modes remain long-standing challenges. Here, we present a detailed fabrication process for scalable vacuum gap capacitors that support ultra-high-coherence mechanical motion, exhibit low microwave loss, and maintain a small footprint compared to planar geometries. We fabricate arrays of up to 24 LC resonators, with capacitors featuring nanometer-scale gap size variations. We demonstrate that the mechanical quality factors can reach up to $40\times 10^{6}$, a 100-fold improvement over other platforms, with microwave quality factors $\mathcal{O}(10^{5})$ at low photon number levels. This platform also achieves a sizable single-photon optomechanical coupling rate of approximately 20Hz. Using this, we cooled the mechanical oscillator to its ground state (0.07 quanta) and squeezed its motion below the vacuum level by 2.7dB [1]. We further demonstrate the scalability of this platform by implementing large-scale optomechanical arrays, a strained graphene model [2], and observing quantum collective phenomena in a mechanical hexamer [3]. These vacuum gap capacitors are promising candidates for coupling superconducting qubits with mechanical systems, serving as storage elements in quantum computing, and exploring gravitational effects on quantum mechanics.
1.   Introduction
Over the past two decades, quantum control of mechanical systems has been firmly established, following the quantum control of individual atoms [4] and ions [5] in the first wave of development, and superconducting circuits [6] in the second wave. This progress has been particularly catalyzed by cavity optomechanics [7], which utilizes radiation-pressure coupling between mechanical oscillators and electromagnetic cavities. More recently, these advancements have been extended to coupling mechanical systems with superconducting qubits via quantum acoustics [6, 8]. These developments have paved the way for quantum optomechanics [9], enabling breakthroughs such as the cooling of low-frequency mechanical oscillators to their quantum ground state [10, 11]—unattainable with passive cooling alone. Other milestones include the generation of entanglement between electromagnetic fields and mechanical oscillators [12, 13], entanglement among macroscopic mechanical oscillators [14, 15], observation of quantum sideband asymmetry in micromechanical oscillators [16, 1], realization of back-action-evading measurements of mechanical motion [17, 18], ponderomotive squeezing of light [19], quantum-coherent coupling of light and mechanical oscillators [20], and real-time quantum feedback control of mechanical oscillators [21]. Furthermore, quantum optomechanics has spurred novel quantum technological innovations, such as interfaces for converting microwave to optical fields with minimal added noise [22, 23] and have been used to amplify microwave signals [24].
A particularly promising platform for the quantum control of mechanical oscillators is circuit optomechanics [25, 10], where a mechanically compliant vacuum-gap capacitor is shunted with an inductor to form a microwave resonator. These capacitors were first introduced in the field of circuit quantum electrodynamics (cQED) to reduce losses in microwave resonators by eliminating the lossy dielectric layer typical of capacitors and increasing the participation ratio of the electric field in vacuum [26]. Over the years, such circuits have enabled remarkable achievements, including mechanical ground-state cooling [10], even below the backaction limit [27], mechanical squeezing [28, 29, 30], entanglement [15, 14, 12, 31], non-classical state storage [32, 33], and non-reciprocal circuits [34, 35, 36]. However, despite these experimental advances, the design and fabrication processes for vacuum-gap capacitors have not kept pace. In particular, circuit optomechanics remains hindered by limited mechanical quality factors, predominantly dictated by the fabrication methods employed for vacuum-gap capacitors, as well as variations in microwave and mechanical properties. These limitations pose significant challenges to scaling up and realizing large-scale lattices.
Figure 1: Overview of the fabrication technique for the next-generation circuit optomechanical platform. a, The main steps of the process consists of etching a trench in the substrate followed by deposition of a sacrificial layer, planarization, top layer definition, release, and finally cool down. Due to the compressive stresses, the top plate will buckle up after the release. However, the drumhead shrinks and flattens at cryogenic temperatures, resulting in a controllable gap size. b, A drumhead parallel plate capacitor after releasing the top layer. c An overview of experimental realizations of circuit optomechanics with vacuum gap capacitors since 2010. The shaded red area highlights the results obtained with our platform, effectively boosting the mechanical coherence by almost two orders of magnitudes.
In this work, we present a comprehensive account of the detailed fabrication process for low-loss vacuum gap capacitors, tailored for circuit quantum optomechanics and circuit quantum electrodynamics applications, and outline the challenges we encountered. We introduce the ”flat-geometry” vacuum gap capacitor, developed to address the limitations of the conventional platform [37, 38, 29, 39], and describe the full optimization process. As shown in Fig. 1a, the flat geometry offers significant advantages, enabling precise control of both the gap and the mechanical frequency through the lithographic process. In Fig. 1b, we provide a SEM micrograph of the top electrode of the vacuum gap capacitor. Furthermore, Fig. 1c illustrates the performance of our fabrication method compared to other approaches reported in the literature. Finally, we demonstrate the application of this device in a circuit optomechanical platform, achieving accurate target frequencies and quality factors up to 40 million. This platform has facilitated the realization of topological lattices in optomechanical circuits [2], the preparation of squeezed mechanical states and the observation of their decoherence in real time [1], and, most recently, collective ground-state cooling [3].
2.   Flat geometry vacuum gap capacitors
The conventional fabrication process of vacuum gap relies on deposition and following lithography definition of a sacrificial layer covering the bottom layer. Following this step, several sacrificial materials on different substrates such as Si3N4 on sapphire [37], polymer on Si [38], SiO2 on quartz [29], and a-Si on sapphire [39] have been tested. In these platforms, due to the deposition-induced compressive stress in the superconducting thin film, the drumhead capacitor buckles up after the release at room temperature, which increases the gap size between two plates up to a few micrometers. Cooling down such devices induces tensile stress in the thin film metal due to the significant difference in thermal expansion rates between the thin film and the substrate. Under tensile stress, the drumhead shrinks, resulting in an approximately 50 nm gap size that is neither predictable nor reproducible. This prevents the precise control of the microwave and mechanical properties of the system at low temperatures and reduces the reproducibility of the design given the high probability of deformations and collapses after the release [40]. In practice, any non-uniformity of the stress distribution in the drumhead after release or asymmetric buckling results in an uncertainty in the final gap size at cryogenic temperatures. This can be in the order of tens of nano-meters, hence limiting the frequency fluctuation in the microwave LC resonator in the order of $\mathcal{O}(10\%)$.
The key idea to overcome the existing challenges, granting better reproducibility and longer mechanical coherence time, is the flat geometry of the vibrating plate. A tensioned vibrating plate results in lower mechanical losses [41] and prevents thermal-induced deformations affecting the capacitor’s gap size. In our fabrication process (Fig. 1a), we first define a trench in a silicon substrate by dry etching. Next, we deposit and pattern the bottom plate of the capacitor inside the trench. The trench is then covered by a thick SiO2 sacrificial layer, which inherits the same topography of the layer underneath. To remove this topography and obtain a flat surface, we use chemical mechanical polishing (CMP) to planarize the SiO2 surface. We then etch back the sacrificial layer down to the substrate layer and deposit the top Al plate of the capacitor. The sacrificial layer will be removed by HF vapor isotropic etching to suspend the structure. After the release process, the drumhead may buckle up (depending on the deposition-induced stress of the thin film) due to the compressive stress; however, at cryogenic temperature the high tensile stress ensures the flatness of the top plate. This will guarantee the gap size to be precisely defined by the trench’s depth and the bottom plate’s thickness. Furthermore, the top plate’s flat geometry significantly reduces the drumhead resonator’s mechanical dissipation. Such advance can be implemented using different materials for substrate, superconducting metal, and sacrificial layer. However, process compatibility of materials and their resilience against different etching steps conduct us to choose a specific set of materials for this process.
[…]
High coherence superconducting circuits are traditionally realized in aluminum [42], niobium [43, 44] and more recently tantalum [45, 46]. With the aim of integrating vacuum gap capacitor in optomechanical system, we decide to use aluminum due to its lighter density which provides larger optomechanical coupling (see Sec. 5 for more details). This makes future integration of our circuits with conventional superconducting qubits more straightforward.
For the sacrificial layer, amorphous silicon (a-Si), silicon nitride (Si3N4), silicon oxide (SiO2), and polymer photoresists are four candidates which were used in the previous generation of circuit optomechanical devices. Each one needs different isotropic etching for the release process. For example, a-Si can be removed by XeF2, which is an exothermic gas etching. Si3N4 can be etched by SF6 plasma, and polymer resists by oxygen plasma. We decided to use SiO2 as our sacrificial layer which can be removed with HF vapor, enabling us to release high aspect ratio structures by avoiding plasma or wet etching, therefore increasing the yield and successful release rate of the process. More details will be discussed in the release section (Sec. 3.5).
3.   Vacuum gap fabrication process
3.1 Etching trenches in silicon
We use optical lithography to transfer patterns on the photoresist. It is performed by direct mask-less optical lithography (${}^{\text{\textregistered}}$ MLA 150). We spin coat a 1 $\mu$m thick ${}^{\text{\textregistered}}$ ECI 3007 positive photoresist after HMDS surface preparation. All photoresist coating and developing steps are processed using automatic coater/developer (${}^{\text{\textregistered}}$ ACS200 GEN3). The exposure dose and depth of focus vary based on the tool and need to be calibrated by dose tests, but are typically set to $\sim~{}150$ mJ/cm2. After the exposure, the resist is developed, and the wafer is rinsed in a spin dryer to clean any unwanted contamination. To remove residual photoresist on the surface of exposed areas, we conduct a short (10-20 seconds) oxygen plasma descum at 200 Watts and 200 sccm (${}^{\text{\textregistered}}$ GiGAbatch). After the descum, the wafer is ready for the etching step.
After lithography, we use deep reactive ion etching (DRIE) to etch the trenches in the silicon substrate. We use $\rm{C_{4}F_{8}}$ chemistry plasma as etchant (${}^{\text{\textregistered}}$ AMS200) with a typical etch rate of $\sim 13$ nm/sec and selectivity of Si:PR10:1:similar-toabsent101$\sim$ 10:1∼ 10 : 1. Due to the small fluctuation of the etch rate in the machine, we use test wafers with a similar pattern to calculate the etch rate by removing the resist and measuring the trench depth using a mechanical profilometer. In addition, we set the total etching $\sim 30$ nm deeper than the target capacitor gap size plus the thickness of the bottom electrode to compensate for potential non-uniformity in the CMP planarization step among different chips on a wafer. The excess depth after CMP can be etched back by IBE in the following steps to reach the desired gap size. The roughness of the silicon inside the trenches is measured $R_{\mathrm{a}}\approx 1.5$ nm with a trench depth uniformity of $\sim 1\%$.
After each etching step, the photoresist is stripped first using UFT remover 1165 wet process, followed by rinse and drying, and then 3 minutes 600 Watt and 400 sccm Oxygen plasma (${}^{\text{\textregistered}}$ GiGAbatch). In the fabrication steps where the wafer contains uncovered thin-film aluminum, it is recommended to reduce either the power or exposure time of the oxygen plasma to avoid additional oxidation and local heating of the metal, specifically for the vibrating top plate.
Figure 2: Etching a trench in silicon. a, Optical microscope image of trenches. b, Atomic Force Microscopy (AFM) of a test trench etched in Si with DRIE. c, AFM of the Si surface inside the trench. The average roughness is $R_{\mathrm{a}}=1.5$ nm. d, An example of a trench etch when the resist descum was not enough. After the etch, the residual photoresist in the trenches results in big hillocks of Si.
3.2 Bottom layer
The first electrode of the vacuum gap capacitor is placed within the silicon trenches. We also add a spiral inductor that will form an LC resonator. For plane circuits, the metal-substrate interface has a major effect on the total loss of the superconducting circuit [47]. While the electric field in the vacuum gap capacitor is mainly stored in the space between the two electrodes, we still clean the wafer after trench etching with Piranha and dip it into HF (1% diluted) for a few minutes to remove the native silicon oxide and minimize as much as possible dielectric loss. Then we rinse and dry the wafer and immediately transfer it to the deposition tool (less than 3 minutes) and pump down the chamber to avoid regrowth of the native oxide.
Deposition of the bottom aluminum layer can be done by either sputtering or electron beam evaporation. However, we find that evaporated films have better thickness uniformity and thickness control compared to sputtered ones. We typically choose 100 nm thickness (Alliance-Concept EVA 760) for the bottom layer.
Figure 3: Bottom Aluminum layer patterning. a, Micrograph of a successfully patterned photoresist inside the trench to etch Al. b and c, Bottom layer circuits after a successful aluminum etching. d (e), Effect of shallow resist on development (etching). The areas close to edges are under exposed. f, When the metal wires pattern are too close to the trench edges, the trench’s edge prevents a proper exposure in corners.
After deposition of the aluminum layer, we repeat the lithography step to pattern the bottom circuit. However, to reduce the topography thickness variation after spin coating deriving from the trenches, we increase the thickness of the resist to 1.2 $\mu$m. In addition, the metal outside of the design area and close to the wafer edge is removed to improve uniformity of the CMP step.
[…]
3.3 Sacrificial layer
Figure 4: SiO2 sacrificial layer deposition. a and b, False color SEM micrograph of the cross section of low temperature oxide low pressure chemical vapor deposition (LTO LPCVD) of SiO2 sacrificial layer covering the trenches. The step coverage and gap filling is perfect, the porosity is low, and the oxide layer is dense. c and d, Plasma enhanced chemical vapor deposition (PECVD) of oxide at 200 Celsius. The oxide layer is porous and does not show a good step coverage, forming void areas at the corners of the trench.
The SiO2 sacrificial layer in our process must satisfy several conditions: 1- It should be grown at low temperatures (below the melting point of aluminum at 660°C) to minimize damage to the aluminum. 2- It should provide good step coverage. 3- It should not be porous, which is important to maintain a flat surface after planarization. 4- It should have high adhesion to the substrate to prevent delamination or dishing when subjected to significant mechanical shear stress during CMP polishing.
PECVD and LTO deposition can both be operated at low temperatures, 100-250°C and 300-450°C, respectively. In Figs. 4c and d, we report SEM image of cross-section of the PECVD silicon oxide deposited at 200 C. The mechanical softness of the oxide is responsible for delamination in CMP, and the porosity produced a large roughness of the aluminum top layer, resulting in lower quality suspended aluminum film. In contrast, in Fig. 4a and b we display an SEM image of LTO-deposited silicon oxide. Such layer presents a higher density and better adhesion to the substrate. Such property are reflected in the CMP step, when the etch/polish rate of LTO-grown oxide is measured $\sim 30\%$ lower than PECVD oxide with same polishing parameters. We find the thickness of the sacrificial layer that optimize the topography removal is around 6 times larger than the maximum topography of the wafer, i.e. the trench depth. For example for a 300 nm trench we deposit 2 $\mu$m oxide layer. This guarantees enough room to run CMP which simultaneously etches and planarizes the surface. We observe less than 0.5% wafer-scale non-uniformity for 3 $\mu$m depositions.
The aluminum film, beneath the sacrificial layer, occasionally displays small and sparse holes (less than 1 $\mu$m diameter) that we attribute to variation in the precursors’ concentration in the LTO chamber, due to previous usage. Nevertheless, we did not observe any sizable impact on the circuits due to this effect.
CMP planarization simultaneously etches the oxide layer and smooths the edges and reduces the topography. For the oxide we deposit, we measure an etching rate between 100 and 300 nm/minute that varies based on pressure, slurry rate and concentration, and rotation speed. A list of the optimal parameters we obtain in our fabrication methods is reported in Table 1.
Table 1: Optimized parameters for one cycle of CMP.
Step Head speed (rpm) Pad speed (rpm) Pressure (Bar) Time (s)
Preparation 40 60 0.2 15
Polishing 78 85 0.4 120
Cleaning 100 100 0.25 30
Slurry 7A5 Gal (pure) or 30N50 (1:1 diluted)
Slurry flow 1/10
Back pressure 0.25 Bar
We find that there is a trade-off between uniformity and residual topography after polishing: longer times reduce the topography (that finally saturates by dishing effect) but increases the thickness non-uniformity at wafer scale. In order to minimize dishing and delamination effect and increase the polishing uniformity, we fill all the empty areas of the wafer between circuits with dummy patterns. These patterns are squares with the size of $60\mu$m with double of this size spacing (see SI for mode details). In Fig. 5, we plot the mechanical profilometer (${}^{\text{\textregistered}}$ Tencor D600) traces across trenches after each of these cycles (with arbitrary translation to make the different profiles comparable). After 10 minutes, all topography is removed and the roughness is smaller than 10 nm as shown in the inset (see SI for more details)
Figure 5: Topography planarization in CMP. CMP enables us to reduce the surface topography from $\sim 500$ nm (the trench depth) to below 10 nm. The figure shows the effect of successive CMP runs on the topography measured by a mechanical profilometer. The final global curve in the topography shows the wafer bow. The inset shows magnified final topography. Adapted from [2, 1].
The thickness of the residual sacrificial layer (inside or outside of the trenches) is measured with an optical spectroscopic reflectometer (${}^{\text{\textregistered}}$ AFT-6100 or ${}^{\text{\textregistered}}$ F54-XY). To measure the thickness inside the trenches, we always locate a ”test trench” with dimension $\sim 0.5\times 0.5$ mm2 (larger than the waist diameter of the optical beam) on every chip to be able to individually measure chips and extract a wafer map of the residual thickness. We aim to have few hundreds of nanometers in thickness for the residual layer above the substrate, as visible in the false color SEM micrograph reported in Fig. 6a. Failing to stop the polishing before this threshold results in large shear stress during CMP that peels off the sacrificial layer from the trench and creates voids around its edge (see Figs. 6b-d). Residual slurry particles (see Figs. 6e and f) are removed by a post-CMP cleaning tool (${}^{\text{\textregistered}}$ Cleaner 428) immediately after the CMP before the wafer dries out.
Figure 6: CMP planarization. a, A cross section SEM showing successful CMP planarization of oxide sacrificial layer covering a trench. The remaining oxide will be removed by IBE to prevent oxide delamination. b and c, SEM of released devices where the CMP reached too close to the substrate surface, resulting in delamination of the sacrificial layer, creating voids at the edges of the trench and creaks on sidewalls of the trenches. d and e, SEM showing the slurry nano-particles after CMP (with PECVD oxide sacrificial layer for these samples). The slurry particles should be cleaned before IBE step using post CMP cleaner mentioned in the text.
The remaining SiO2 on the Si surface is etched by argon milling (${}^{\text{\textregistered}}$ Nexus IBE350) with a slow etch rate of 35 nm/minute and 1:1 etch selectivity for Si:SiO2, which increases the controllability of the residual layer thickness and avoid increasing topography when the etching transitions from SiO2 to Si. To further improve uniformity, the wafer rotates during etching at 10 rpm. In addition, to avoid redepositions, the wafer is tilted by 45 during etching. Post-etching surface roughness measurements on Si yields to $R_{\mathrm{a}}\simeq 0.75$ nm demonstrating negligible surface damage. We target an over-etch of $\sim 20$ nm to make sure all SiO2 is removed from surface. See Fig. 7 to see the oxide in the trench after IBE.
Figure 7: IBE etch-back. a and b, SEM images showing trenches after planarization and IBE etch-back. The oxide-silicon border is dense and smooth, making a perfect condition for top-layer deposition.
To have electrical access to the bottom electrode, we open a via through the silicon oxide. We define a square pad of $\sim 20\mu\rm{m}\times 20\mu\rm{m}$ (in the trench) galvanically connected to the bottom plate of the capacitor. With the same lithography procedure described in the beginning, we pattern a smaller rectangle on the resist on top of the SiO2 layer which covers the bottom connection pad. To have a smooth metal coverage, we reflow the resist by heating up the wafer to 180 Celsius for 30 seconds using a standard hot plate. After the reflow, we do the standard descum to remove resist residues. Afterward, we use DRIE plasma etching (${}^{\text{\textregistered}}$ APS) with $\rm{CHF_{3}}$ chemistry which offers 1:1 selectivity for $\rm{SiO_{2}}$:photoresist and transfer the photoresist pattern into the oxide. Then the resist is removed by the standard procedure discussed earlier. Avoiding reflow produces thin aluminum contact ($<50$ nm) on the edges of the galvanic connection as displayed in Figs. 8a and b. In this case, we observe strong high-power nonlinearities in the microwave response, that we attribute to high local current densities in the connection region. The galvanic connections after the reflow step, reported in Figs. 8c and d, do not show non-idealities.
Figure 8: Problem of galvanic connection with sharp edges. a, The galvanic connection on SiO2 openings with sharp edges. Due to the local thickness decrease of aluminum layer on the sharp edges, these circuits show frequency shift when the intra-cavity photon number -in other words, circulating current- is high. b, In addition, the sharp edges of the opening may result in accumulation of aluminum during the top layer deposition. c and d, SEM of a galvanic connection with resist reflow process after the release showing smooth transition of the top layer to the lower level.
3.4 Top aluminum layer deposition
We tested two different electron beam evaporators (Eva: ${}^{\text{\textregistered}}$ EVA 760, and Plassys: ${}^{\text{\textregistered}}$ MEB550SL3) for the deposition of the top aluminum layer. Eva is a standard electron beam evaporator with 450 mm working distance. A 200 nm aluminum film grown with 0.5 nm/s deposition rate in Eva results in $\sim 50$ MPa compressive stress, that produces a buckling of the drumhead in a dome shape after the release at room temperature. This effect has the advantage of improving the release yield since the gap size will increase more than $\sim 1\mu$m, and HF vapor can penetrate easier to remove the remaining sacrificial layer. With this method, however, a thin layer of native aluminum oxide remains in the top layer to the bottom layer connection (through the pad that was explained in the previous section), making it a resistive connection.
Figure 9: Released devices. a to c, SEM image, microscope image, and optical profilometry of released drumhead capacitors for the case of near-zero or tensile stress (Plassys deposition), receptively. d to f, SEM image, microscope image, and optical profilometry of released drumhead capacitors for the case of compressive stress in the top aluminum layer (using Eva evaporation machine). Drums with compressive stress buckle up to $\sim 2\mu$m and form a dome shape visible under the microscope. However the gap size at cryogenic temperatures goes to the designed value due to the temperature induced tensile stress and consequently flatness of the drum.
To overcome this drawback, we argon-ion mill and deposit the metal without breaking the vacuum in Plassys. This can help to significantly reduce the microwave cavity heating when sending high power to the device [1]. An optimized evaporation on 200 nm Al in Plassys gives minor tensile stress in the film. Finally, an Al2O3 layer is grown on the drumhead resonator by injecting 10 mBar of 99.99% pure oxygen in Plassys oxidation chamber. Additional deposition parameters are discussed in Sec. 5.
After the deposition, we use the standard lithography technique to pattern the top layer and etch it using the same wet process which has been mentioned earlier.
3.5 Release
We dice the wafer into chips before the release (see details on dicing in SI). The last step of the fabrication process is releasing the vacuum gap capacitors by removing all $\rm{SiO_{2}}$ sacrificial layer from trenches. This is done by vapor phase Hydrofluoric acid etching (${}^{\text{\textregistered}}$ uEtch). Reaction with the sacrificial SiO2 on the wafer surface produces silicon tetrafluoride (SiF4) gas and water vapor:
$\rm{SiO_{2}}+\rm{4HF}^{-2}+\rm{4C_{2}H_{5}OH}^{+2}\longrightarrow\rm{SiF_{4}}+\rm{2H_{2}O}+\rm{4C_{2}H_{5}OH}$ (1)
Although the liquid HF attacks aluminum, we observe that the vapor HF does not deteriorate or degrade Al films. We use a recipe with 125 Torr pressure and $\sim 100$ nm/minute etch rate for 900 seconds etching time in every cycle. Since the vapor HF needs to penetrate horizontally between the top Al layer and the trench bottom surface, the total number of cycles needed should be calculated based on the maximum lateral distance between two penetration windows for the gas etchant, considering the pattern of Al covering the trenches. We multiply this number by a factor of four to prevent any residual oxide and to ensure the whole structure is released. The etching process is liquid-free, which is crucial to release structures with high aspect ratios, in our case $\sim 100$ $\mu$m big drums suspended by $\sim 200$ nm gap above another metallic layer. Any water formation will result in the collapse and sticking of two capacitor plates.
Examples of successfully released devices are shown in Figs. 9 and 10.
Figure 10: Elements of the final device. a, Focused ion beam cross-section of a capacitor before removing the SiO2 sacrificial layer (Pt is used as the focused ion beam protective layer). The flatness of the top layer is visible in the image indicating a successful CMP planarization. b, SEM image of the perforated released drumhead. The holes facilitate the release process. c, Magnified SEM of one release hole. The bottom Al layer is visible from the hole. The small particles seen on the aluminum surface are aluminum hillocks, a well-know accumulation of aluminum in evaporation technique. The size and distribution of hillocks depends on the evaporation rate and the pressure of the chamber. d, The bottom wire going under a drum. e, A magnified SEM of a drumhead clamp. f and g, The spiral inductor air bridges and the galvanic connection.
To facilitate the release, specifically for big drums, we perforate drums by defining small holes with 1.8 $\mu$m diameter and distance of $\sim 10\mu$m as shown in Fig. S14.
We successfully release vacuum gaps with the gap size down to 75 nm, but the success rate of the release was not high for such a low gap size (Fig. S13). For gaps above 150 nm, we find almost 100% yield of release. The success rate depends on the trench size, the thickness of the top plate, and the room temperature stress. Smaller drums have a lower risk of collapse and can tolerate smaller gap sizes. We usually use 150-200 nm thick Al top layer. Attempts to release thin top layers (50 nm) were not successful (see Fig. S15).
4.   Measurement and results
To determine the mechanical and microwave properties of the vacuum gap capacitor, we shunt it by a meander inductor and thermally anchor the sample to the mixing chamber of a dilution refrigerator at 10 mK. To determine the mechanical frequency of the vacuum gap capacitor we use optomechanically induced transparency (OMIT) [48].
The bare mechanical damping rates can be either characterized by extracting the linewidth of the OMIT features while pumping the microwave resonance with low powers corresponding to optomechanical cooperativities below one ($\mathcal{C}\ll 1$) or conducting a ring-down experiment. Since the presented fabrication process results in sub-Hertz damping rates of mechanical modes, the OMIT measurement requires low frequency resolution and typically slow sweeping speed. Therefore, the ringdown measurement is more suitable for such characterizations. For higher red-detuned probe powers the effective mechanical damping rate is $\Gamma_{\mathrm{tot}}=\Gamma_{\mathrm{m}}(1+\mathcal{C})$. Sweeping the power of the red probe enables us to directly measure $\Gamma_{\mathrm{m}}$ (Figs. 11c and d).
Figure 11: Ringdown measurement of a high-$Q_{\mathrm{m}}$ electromechanical system. a, Pulse and frequency scheme of the ringdown measurement. A strong blue detuned pulse is exciting the mechanical oscillator through optomechanical parametric instability. A red-detuned readout pulse generates an optomechanical sideband on resonance. b, Example of ringdown traces measured for different readout powers. The initial nonlinear behavior in the ringdown trace may be due to the energy exchange between different mechanical modes of the drumhead at high amplitude vibrations. For the exponential fitting (shown by green solid lines), we only use the low-power linear part of the ring down. The trace corresponding to pump power of -50 dBm shows $Q_{\mathrm{m}}=4\times 10^{7}$. Adapted from [1].
We studied 16 separate electromechanical LC circuits fabricated on a 9.5 mm×$\times$6.5 mm chip (See SI for the chip design). The frequencies of microwave and mechanical resonances are multiplexed in the chip in the range of 5-7 GHz and 1.5-2.5 MHz, respectively. This is done by changing the trench radius for mechanical frequency tuning (from 60 $\mu$m to 100 $\mu$m), and the capacitor bottom plate radius for microwave frequency tuning. All 16 LC circuits were magnetically coupled to a micro-strip waveguide.
We provide the measured quality factors and mechanical frequencies for a chip with 16 independent electromechanical LC resonators in a chip - we did not observe two LC resonators, due to overlapping their frequencies with the stop-band of a Josephson traveling wave parametric amplifier [49] used in the measurement chain. More than 50% of the resonators exhibit above $20\times 10^{6}$ mechanical quality factor (Fig. 13a), which demonstrates a high yield in our new fabrication process.
A persistent challenge in circuit optomechanical platforms has been the heating of the microwave cavity caused by high circulating currents in the circuit [10], which limits the minimum occupation of the mechanical oscillator using sideband cooling. As discussed in Sec. 3.4, argon milling of the bottom-layer aluminum prior to making the galvanic connection removes the resistive aluminum oxide layer, thereby reducing microwave loss. Figure 12 shows the microwave cavity heating with and without ion milling treatment, demonstrating that this method reduces heating by an order of magnitude. With this improvement, the microwave resonator quality factor is limited by dielectric loss of its capacitive element. By measuring the internal microwave quality factor as a function of the intracavity photon number, we report internal quality factors up to $10^{5}$ for photon number $n\sim\mathcal{O}(100)$ going up to $2\times 10^{6}$ for photon number $n\sim\mathcal{O}(10^{7})$ for a resonance frequency $\sim\mathcal{O}(5\,\rm{GHz})$.
Figure 12: Microwave cavity heating treatment. Heating of the microwave cavity (in units of quanta) as a function of the red pump cooperativity for a device without ion milling treatment (red circles) and with milling treatment (blue circles). Adapted from [1].
Figure 13: Quality factor distribution and Extraction of the aluminum film stress at low temperatures. a, Measured mechanical quality factors versus the mechanical frequencies in a chip with 14 separate electromechanical LC circuits. Outlier quality factors may be caused by post-fabrication device contamination. b, Measured mechanical frequencies on a chip with 16 separate electromechanical LC resonators. In the chip layout the trench radius of drums is swept from 60 $\mu$m to 100 $\mu$m. Since two LC resonances were inaccessible due to frequency overlap with the JTWPA stopband, we measured 14 resonances and considered an error bar showing the uncertainty in the trench radii. The red line shows the theoretical curve with $\sigma_{\text{Al}}=350$ MPa and the green shade shows theory bounds corresponding to $350\text{MPa}\times(1\pm 10\%)$ uncertainty.
The fundamental mechanical frequency of an ideal and fully clamped drum is given by:
$\Omega_{\mathrm{m}}=\frac{\alpha_{0,1}}{R}\sqrt{\frac{\sigma_{\mathrm{m}}}{\rho}},$ (2)
where $\alpha_{0,1}$ is the first root of the zeroth order Bessel function ($J_{0}$), and $\sigma_{\mathrm{m}}$ and $\rho$ are the mechanical stress and density of the material, respectively, in our case aluminum with $\rho_{\mathrm{Al}}=2700\ \mathrm{kg}/\mathrm{m}^{3}$.
To experimentally extract the value of the mechanical stress in drumheads at low temperature, we plot mechanical frequencies versus trench radius and fit the theoretically expected frequencies to extract the stress as $\sigma_{\text{Al}}=350(\pm 10\%)$ MPa, as shown in Fig. 13b. Tapering the clamping points will increase the local stress proportional to the tapering ratio [41]. We swept the clamp ratio of drums with 50 $\mu$m radius and observed broken legs for the ones with more than $\sim 1$ GPa stress. This can indicate an upper limit for the yield stress of the aluminum thin films at low temperatures which has not been reported in the literature. Detailed discussion on the aluminum thin film stress at low temperatures is provided in SI.
The quality factor of a nano-mechanical oscillator is expressed as $Q_{\rm m}=Q_{0}\times D_{Q}$, where $Q_{0}$ represents the material quality factor, and $D_{Q}$ is the loss dilution factor [50, 51]. The dilution factor $D_{Q}$ depends on the oscillator’s geometry and material properties such as stress and Young’s modulus. Using finite element simulations of our drum resonator in COMSOL (Fig. 14), we estimate the dilution factor in our device to be $D_{Q}\approx 100$. Additionally, the aluminum quality factor at 10 mK is estimated to be $Q_{0}\approx 4\times 10^{5}$.
Figure 14: Dilution factor estimation. Mechanical dilution factor as a function of the drum resonator radius, obtained through finite element method simulations in COMSOL. The inset illustrates the simulated drum model in COMSOL. Adapted from [1].
Optomechanical building blocks can be couple with each other to form arrays and lattices [52, 53, 54, 55, 56, 57]. Namely, such multimode systems can implement Su-Schrieffer-Heeger (SSH) arrays [58, 59], as a fundamental topological model. Such optomechanical arrays can exhibit non-trivial topological properties, localized edge states, and be exploited to directly measure hybridized microwave modeshapes to fully reconstruct the system’s Hamiltonian [2]. To demonstrate the scalability of our process, we fabricated a 12-site 1D array of identical electromechanical LC circuits with alternating mutual inductance realizing SSH model (Fig. 15a). We slightly swept the trench radius of drumhead resonators in the array (by 500 nm) to shift their frequencies and be able to identify them in the OMIT response (Fig. 15b). As shown in Fig. 15c, the measured mechanical frequencies from OMIT experiment perfectly follow the theoretical fit. The standard deviation of mechanical frequency disorder from the theoretical fit is less than 0.2%, which shows a perfect control and high reproducibility in the process.
Figure 15: Low-disorder optomechanical arrays. a, a 12-site electromechanical array realizing SSH model. All sites are identical, but with slightly different trench radius (between 50 $\mu$m to 55.5 $\mu$m with 0.5 $\mu$m increment) resulting in mechanical frequency. b, OMIT response showing all 12 mechanical modes. The normalized scattering parameter is shown versus the pump-probe frequency detuning. c, Measured mechanical frequencies versus the designed trench radius (dots). The red line shows theoretical fit with $1/R$ scaling rule.
5.   Discussion and Conclusion
In summary, we report the nano-fabrication technique for ultra-coherent and reproducible superconducting circuit optomechanics. Such systems have shown millisecond scale quantum decoherence [1], were used to make large-scale optomechanical lattices [2], and can be used to demonstrate quantum collective dynamics of mechanical oscillators [3].
Having such controllable vacuum gap capacitor allows to realize resonators with precise resonance frequencies [3], on par with other platform as kinetic inductance resonators [60] or Josephson junction-based circuitry [61].
The presented ultra-coherent electromechanical system can be exploited in quantum sensing applications [62]. The expected on-resonance force sensitivity of our device can be estimated as $\sqrt{S_{FF}}=\sqrt{2k_{\mathrm{B}}Tm_{\mathrm{eff}}\Gamma_{\mathrm{m}}}\simeq 200\times 10^{-21}\left[\frac{\text{N}}{\sqrt{\mathrm{Hz}}}\right]$, which is considerably low compared with several other optomechanical platforms thanks to the high mechanical quality factor and the low operating temperature of the device ($T=10$ mK and $m_{\mathrm{eff}}\simeq 2$ ng). Moreover, such high-Q electromechanical system may benefit the implementation of qubit-mechanics interfaces [32], generation of mechanical non-classical states [63], realization of long life-time memories for quantum computation and communication [64, 65], and it may set the stage to perform fundamental tests of quantum mechanics in macroscopic scales such as quantum gravity tests [66, 67], high fidelity Bell tests [68, 69], quantum teleportation [70], or even the search for Dark matter [71, 72].
Acknowledgement
This work was supported by the Swiss National Science Foundation (SNSF) under grant No. 231403 (CoolMe). M.S. acknowledges support from the EPFL Center for Quantum Science and Engineering postdoctoral fellowship. All devices were fabricated in the Center of MicroNanoTechnology (CMi) at EPFL.

### v2

@@
 Abstract
-Vacuum gap capacitors have recently gained considerable attention in superconducting circuit platforms due to their compact design and low dielectric losses in the microwave regime.
-Their ability to support mechanical vibrational modes makes them ideal candidates for circuit optomechanics.
-However, precise control of gap size and achieving high coherence in mechanical modes remain long-standing challenges.
-Here, we present a detailed fabrication process for scalable vacuum gap capacitors that support ultra-high-coherence mechanical motion, exhibit low microwave loss, and maintain a small footprint compared to planar geometries.
-We fabricate arrays of up to 24 LC resonators, with capacitors featuring nanometer-scale gap size variations.
-We demonstrate that the mechanical quality factors can reach up to $40\times 10^{6}$, a 100-fold improvement over other platforms, with microwave quality factors $\mathcal{O}(10^{5})$ at low photon number levels.
-This platform also achieves a sizable single-photon optomechanical coupling rate of approximately 20Hz.
-Using this, we cooled the mechanical oscillator to its ground state (0.07 quanta) and squeezed its motion below the vacuum level by 2.7dB [1].
-We further demonstrate the scalability of this platform by implementing large-scale optomechanical arrays, a strained graphene model [2], and observing quantum collective phenomena in a mechanical hexamer [3].
-These vacuum gap capacitors are promising candidates for coupling superconducting qubits with mechanical systems, serving as storage elements in quantum computing, and exploring gravitational effects on quantum mechanics.
+Vacuum-gap capacitors have recently attracted significant interest in superconducting circuit platforms due to their compact design and exceptionally low dielectric losses in the microwave regime.
+Their intrinsic ability to support mechanical vibrational modes makes them well-suited for circuit optomechanics.
+However, precise control over the gap size and the realization of high-coherence mechanical modes remain longstanding challenges.
+Here, we present a detailed and scalable fabrication process for vacuum-gap capacitors that support ultra-high-coherence mechanical motion, exhibit low microwave loss, and occupy a significantly smaller footprint compared to conventional planar geometries.
+By employing a planarized SiO2 sacrificial layer, we achieve vacuum gaps on the order of 150 nm.
+Using this platform, we have recently demonstrated ground-state cooling and motion squeezing of a mechanical oscillator with a quality factor of 40 million—a 100-fold improvement compared to prior works—as well as a single-photon optomechanical coupling rate of approximately 15Hz [1].
+Additional achievements include the realization of an optomechanical topological lattice with 24 sites [2] and the observation of quantum collective dynamics in a mechanical hexamer [3].
+Collectively, these results underscore the potential of vacuum-gap capacitors as a platform for coupling superconducting qubits to mechanical systems, enabling quantum storage, and probing gravitational effects in quantum mechanics.
 1.   Introduction
@@
 In this work, we present a comprehensive account of the detailed fabrication process for low-loss vacuum gap capacitors, tailored for circuit quantum optomechanics and circuit quantum electrodynamics applications, and outline the challenges we encountered.
-We introduce the ”flat-geometry” vacuum gap capacitor, developed to address the limitations of the conventional platform [37, 38, 29, 39], and describe the full optimization process.
+We introduce the “flat-geometry” vacuum gap capacitor, developed to address the limitations of the conventional platform [37, 38, 29, 39], and describe the full optimization process.
 As shown in Fig.
@@
 Furthermore, Fig.
-1c illustrates the performance of our fabrication method compared to other approaches reported in the literature.
-Finally, we demonstrate the application of this device in a circuit optomechanical platform, achieving accurate target frequencies and quality factors up to 40 million.
+1c illustrates the performance of our fabrication method compared to other approaches reported in the literature, excluding metalized membranes that feature very large footprint [40, 41] (see supplementary information).
+Finally, we demonstrate the application of this device in a circuit optomechanical platform, achieving accurate target frequencies and quality factors as high as 40 million.
 This platform has facilitated the realization of topological lattices in optomechanical circuits [2], the preparation of squeezed mechanical states and the observation of their decoherence in real time [1], and, most recently, collective ground-state cooling [3].
@@
 The conventional fabrication process of vacuum gap relies on deposition and following lithography definition of a sacrificial layer covering the bottom layer.
-Following this step, several sacrificial materials on different substrates such as Si3N4 on sapphire [37], polymer on Si [38], SiO2 on quartz [29], and a-Si on sapphire [39] have been tested.
+Several sacrificial materials on different substrates such as Si3N4 on sapphire [37], polymer on Si [38], SiO2 on quartz [29], and a-Si on sapphire [39] have been tested.
 In these platforms, due to the deposition-induced compressive stress in the superconducting thin film, the drumhead capacitor buckles up after the release at room temperature, which increases the gap size between two plates up to a few micrometers.
@@
 High coherence superconducting circuits are traditionally realized in aluminum [44], niobium [45, 46] and more recently tantalum [47, 48].
-With the aim of integrating vacuum gap capacitor in optomechanical system, we decide to use aluminum due to its lighter density which provides larger optomechanical coupling (see Sec.
-5 for more details).
+With the aim of integrating vacuum gap capacitor in optomechanical system, we decide to use aluminum due to its lighter density which provides larger optomechanical coupling (see SI for more details).
 This makes future integration of our circuits with conventional superconducting qubits more straightforward.
@@
 We decided to use SiO2 as our sacrificial layer which can be removed with HF vapor, enabling us to release high aspect ratio structures by avoiding plasma or wet etching, therefore increasing the yield and successful release rate of the process.
-More details will be discussed in the release section (Sec.
+In addition, it has infinite selectivity to aluminum and silicon.
+More details are given in the release section (Sec.
 3.5).
@@
 All photoresist coating and developing steps are processed using automatic coater/developer (${}^{\text{\textregistered}}$ ACS200 GEN3).
-The exposure dose and depth of focus vary based on the tool and need to be calibrated by dose tests, but are typically set to $\sim~{}150$ mJ/cm2.
+The exposure dose and depth of focus vary based on the tool and need to be calibrated by dose tests, but are typically set to $\sim~{}150$ mJ/cm2 and 0, respectively.
 After the exposure, the resist is developed, and the wafer is rinsed in a spin dryer to clean any unwanted contamination.
@@
 The roughness of the silicon inside the trenches is measured $R_{\mathrm{a}}\approx 1.5$ nm with a trench depth uniformity of $\sim 1\%$.
-After each etching step, the photoresist is stripped first using UFT remover 1165 wet process, followed by rinse and drying, and then 3 minutes 600 Watt and 400 sccm Oxygen plasma (${}^{\text{\textregistered}}$ GiGAbatch).
+After each etching step, the photoresist is stripped first using UFT remover 1165 wet process, followed by rinse and drying, and then 3 minutes 200 Watt and 200 sccm Oxygen plasma (${}^{\text{\textregistered}}$ GiGAbatch).
 In the fabrication steps where the wafer contains uncovered thin-film aluminum, it is recommended to reduce either the power or exposure time of the oxygen plasma to avoid additional oxidation and local heating of the metal, specifically for the vibrating top plate.
@@
 We also add a spiral inductor that will form an LC resonator.
-For plane circuits, the metal-substrate interface has a major effect on the total loss of the superconducting circuit [47].
+For microwave circuits, the metal-substrate interface has a major effect on the total loss of the superconducting circuit [49].
 While the electric field in the vacuum gap capacitor is mainly stored in the space between the two electrodes, we still clean the wafer after trench etching with Piranha and dip it into HF (1% diluted) for a few minutes to remove the native silicon oxide and minimize as much as possible dielectric loss.
@@
 However, we find that evaporated films have better thickness uniformity and thickness control compared to sputtered ones.
-We typically choose 100 nm thickness (Alliance-Concept EVA 760) for the bottom layer.
+We typically choose 100 nm thickness (Alliance-Concept EVA 760) for the bottom layer, deposited with 0.5 nm/sec rate.
 Figure 3: Bottom Aluminum layer patterning.
@@
 The step coverage and gap filling is perfect, the porosity is low, and the oxide layer is dense.
-c and d, Plasma enhanced chemical vapor deposition (PECVD) of oxide at 200 Celsius.
+c and d, Plasma enhanced chemical vapor deposition (PECVD) of oxide at ${200^{\circ}}{\rm{C}}$.
 The oxide layer is porous and does not show a good step coverage, forming void areas at the corners of the trench.
@@
 In order to minimize dishing and delamination effect and increase the polishing uniformity, we fill all the empty areas of the wafer between circuits with dummy patterns.
-These patterns are squares with the size of $60\mu$m with double of this size spacing (see SI for mode details).
+These patterns are squares with the size of $60\mu$m with double of this size spacing (see SI for more details).
 In Fig.
 5, we plot the mechanical profilometer (${}^{\text{\textregistered}}$ Tencor D600) traces across trenches after each of these cycles (with arbitrary translation to make the different profiles comparable).
-After 10 minutes, all topography is removed and the roughness is smaller than 10 nm as shown in the inset (see SI for more details)
+After 10 minutes, topography is removed to below 10 nm, as it is shown in the inset (see SI for more details).
 Figure 5: Topography planarization in CMP.
@@
 Finally, an Al2O3 layer is grown on the drumhead resonator by injecting 10 mBar of 99.99% pure oxygen in Plassys oxidation chamber.
-Additional deposition parameters are discussed in Sec.
-5.
+Additional deposition parameters are discussed in supplementary information (SI).
 After the deposition, we use the standard lithography technique to pattern the top layer and etch it using the same wet process which has been mentioned earlier.
@@
 We dice the wafer into chips before the release (see details on dicing in SI).
-The last step of the fabrication process is releasing the vacuum gap capacitors by removing all $\rm{SiO_{2}}$ sacrificial layer from trenches.
+The last step of the fabrication process is releasing the vacuum gap capacitors by removing the $\rm{SiO_{2}}$ sacrificial layer.
 This is done by vapor phase Hydrofluoric acid etching (${}^{\text{\textregistered}}$ uEtch).
-Reaction with the sacrificial SiO2 on the wafer surface produces silicon tetrafluoride (SiF4) gas and water vapor:
+Reaction with the sacrificial SiO2 on the wafer surface (in the presence of ethanol as the catalyst) produces silicon tetrafluoride (SiF4) gas and water vapor:
 $\rm{SiO_{2}}+\rm{4HF}^{-2}+\rm{4C_{2}H_{5}OH}^{+2}\longrightarrow\rm{SiF_{4}}+\rm{2H_{2}O}+\rm{4C_{2}H_{5}OH}$ (1)
@@
 The etching process is liquid-free, which is crucial to release structures with high aspect ratios, in our case $\sim 100$ $\mu$m big drums suspended by $\sim 200$ nm gap above another metallic layer.
-Any water formation will result in the collapse and sticking of two capacitor plates.
-Examples of successfully released devices are shown in Figs. 9 and 10.
+Any liquid formation will result in the collapse and sticking of two capacitor plates.
+Examples of successfully released devices are shown in Figs.
+9 and 10.
 Figure 10: Elements of the final device.
-a, Focused ion beam cross-section of a capacitor before removing the SiO2 sacrificial layer (Pt is used as the focused ion beam protective layer).
+a, SEM micrograph of a focused ion beam cross-section of a capacitor before removing the SiO2 sacrificial layer (Pt is used as the focused ion beam protective layer).
 The flatness of the top layer is visible in the image indicating a successful CMP planarization.
@@
 f and g, The spiral inductor air bridges and the galvanic connection.
-To facilitate the release, specifically for big drums, we perforate drums by defining small holes with 1.8 $\mu$m diameter and distance of $\sim 10\mu$m as shown in Fig. S14.
-We successfully release vacuum gaps with the gap size down to 75 nm, but the success rate of the release was not high for such a low gap size (Fig.
-S13).
+To facilitate the release, specifically for big drums, we perforate drums by defining small holes with 1.8 $\mu$m diameter and distance of $\sim 10\mu$m as shown in Fig.
+10.
+We successfully release vacuum gaps with the gap size down to 75 nm, but the success rate of the release was not high for such a low gap size (see SI).
 For gaps above 150 nm, we find almost 100% yield of release.
@@
 We usually use 150-200 nm thick Al top layer.
-Attempts to release thin top layers (50 nm) were not successful (see Fig.
-S15).
+Attempts to release thin top layers (50 nm) were not successful (see SI).
 4.   Measurement and results
@@
 The bare mechanical damping rates can be either characterized by extracting the linewidth of the OMIT features while pumping the microwave resonance with low powers corresponding to optomechanical cooperativities below one ($\mathcal{C}\ll 1$) or conducting a ring-down experiment.
-Since the presented fabrication process results in sub-Hertz damping rates of mechanical modes, the OMIT measurement requires low frequency resolution and typically slow sweeping speed.
+Since the presented fabrication process results in sub-Hertz damping rates of mechanical modes, the OMIT measurement requires low frequency resolution and typically low resolution bandwidth.
 Therefore, the ringdown measurement is more suitable for such characterizations.
-For higher red-detuned probe powers the effective mechanical damping rate is $\Gamma_{\mathrm{tot}}=\Gamma_{\mathrm{m}}(1+\mathcal{C})$.
+For higher red-detuned probe powers, the effective mechanical damping rate is $\Gamma_{\mathrm{tot}}=\Gamma_{\mathrm{m}}(1+\mathcal{C})$.
 Sweeping the power of the red probe enables us to directly measure $\Gamma_{\mathrm{m}}$ (Figs.
@@
 13a), which demonstrates a high yield in our new fabrication process.
-A persistent challenge in circuit optomechanical platforms has been the heating of the microwave cavity caused by high circulating currents in the circuit [10], which limits the minimum occupation of the mechanical oscillator using sideband cooling.
+In Fig.
+13a, we reported the mechanical quality factors of 14 devices fabricated on a single chip.
+In Fig.
+14, we present data from a much broader dataset comprising 119 mechanical oscillators measured over a span of three years, encompassing three different fabrication batches across six wafers.
+The fabrication of batches CCv3, CCv4, and CSv1 was completed in September 2022, May 2023, and September 2024, respectively.
+The average mechanical quality factor across 119 devices is remarkably high, at 4.3 million.
+These results demonstrate the high yield and consistency of our fabrication process across batches and over time.
+A persistent challenge in circuit optomechanical platforms has been the heating of the microwave cavity caused mainly by high circulating currents in the circuit [10], which limits the minimum occupation of the mechanical oscillator using sideband cooling.
 As discussed in Sec.
@@
 With this improvement, the microwave resonator quality factor is limited by dielectric loss of its capacitive element.
-By measuring the internal microwave quality factor as a function of the intracavity photon number, we report internal quality factors up to $10^{5}$ for photon number $n\sim\mathcal{O}(100)$ going up to $2\times 10^{6}$ for photon number $n\sim\mathcal{O}(10^{7})$ for a resonance frequency $\sim\mathcal{O}(5\,\rm{GHz})$.
+By measuring the internal microwave quality factor as a function of the intracavity photon number, in [3] we report internal quality factors up to $10^{5}$ for photon number $n\sim\mathcal{O}(100)$ going up to $2\times 10^{6}$ for photon number $n\sim\mathcal{O}(10^{7})$ for a resonance frequency $\sim\mathcal{O}(5\,\rm{GHz})$.
+The reported microwave quality factor is compatible with the large participation ratio of the electric field in the native aluminum oxide present on the capacitor leads and other vacuum gap capacitors have shown similar performance in compact capacitor for resonator [52] and for transmon qubit [53].
 Figure 12: Microwave cavity heating treatment.
@@
 The red line shows the theoretical curve with $\sigma_{\text{Al}}=350$ MPa and the green shade shows theory bounds corresponding to $350\text{MPa}\times(1\pm 10\%)$ uncertainty.
+Figure 14: Batch-to-batch mechanical quality factors.
+The mechanical quality factors for three different batches CCv3, CCv4, and CSv1, which are fabricated on September 2022, May 2023, and September 2024, respectively.
+The transparent region is the estimated density of quality factors, the top line show the maximum quality factor, the middle line shows the average quality factor, and the bottom line shows the minimum quality factor.
+All the drums have 70 $\mu$m radius (average mechanical frequency 2 MHz).
 The fundamental mechanical frequency of an ideal and fully clamped drum is given by:
@@
 Using finite element simulations of our drum resonator in COMSOL (Fig.
-14), we estimate the dilution factor in our device to be $D_{Q}\approx 100$.
+15), we estimate the dilution factor in our device to be $D_{Q}\approx 100$.
 Additionally, the aluminum quality factor at 10 mK is estimated to be $Q_{0}\approx 4\times 10^{5}$.
@@
 Adapted from [1].
-Optomechanical building blocks can be couple with each other to form arrays and lattices [52, 53, 54, 55, 56, 57].
+Optomechanical building blocks can be coupled with each other to form arrays and lattices [56, 57, 58, 59, 60, 61].
 Namely, such multimode systems can implement Su-Schrieffer-Heeger (SSH) arrays [62, 63], as a fundamental topological model.
@@
 To demonstrate the scalability of our process, we fabricated a 12-site 1D array of identical electromechanical LC circuits with alternating mutual inductance realizing SSH model (Fig.
-15a).
-We slightly swept the trench radius of drumhead resonators in the array (by 500 nm) to shift their frequencies and be able to identify them in the OMIT response (Fig.
-15b).
+16a).
+We slightly sweep the trench radius of drumhead resonators in the array (by 500 nm) to shift their frequencies and be able to identify them in the OMIT response (Fig.
+16b).
 As shown in Fig.
-15c, the measured mechanical frequencies from OMIT experiment perfectly follow the theoretical fit.
+16c, the measured mechanical frequencies from OMIT experiment perfectly follow the theoretical fit.
 The standard deviation of mechanical frequency disorder from the theoretical fit is less than 0.2%, which shows a perfect control and high reproducibility in the process.
@@
 All devices were fabricated in the Center of MicroNanoTechnology (CMi) at EPFL.
+Data availability
+The data that support the findings of this article are openly available [77].